Please help solve this problem
first day 5 adult tickets and 8 children tickets were sold=68
second day 3 adult and 5 childrens=42. how much are the adult tickets

Answer by stanbon(75887)

first day 5 adult tickets and 8 children tickets were sold=68
second day 3 adult and 5 childrens=42. how much are the adult tickets
----------------
5a + 8c = 68
3a + 5c = 42
---------------------
Multiply 1st by 5
Multiply 2nd by 8
---------------------
25a + 40c = 5*68
24a + 40c = 8*42
---------------------
Subtract:
a = 340-336
a = 4 (# of adult tickets sold)
-----------
Cheers,
Stan H.
============